# Modeling Developer Burnout with GenAI Adoption

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Empirical studies in collaborative and social computingccs: Human-centered computing
@@ 1. Introduction
We are now at a pivotal shift in how Generative AI (GenAI) is transforming how we work and live. Since late 2022, GenAI has progressed from experimental pilots to mainstream adoption, reshaping how code is written, reviewed, and maintained (Nguyen-Duc et al., 2025; Oliver Wyman Forum, 2024; Maslej et al., 2025). Individual adoption is rapid—84% of developers—in the 2025 Stack Overflow survey reported using or planning to use GenAI tools (Stack Overflow, 2025b). Organizations are not far behind. Motivated by promises of higher productivity, they are heavily investing in GenAI (GitHub, 2023; Bick et al., 2025).
@@
Some studies support these investments. For example, GitHub reported that developers completed coding tasks 55.8% faster (GitHub, 2025). Yet, others report increased costs. For instance, the 2025 State of Software Delivery Report by Harness (Harness, 2025) reported 67% of developers spent more time debugging AI-generated code and 68% spent more time fixing AI-created security issues. Similarly, Becker et al. (2025) reported an overall 19% productivity loss in their observation study. These contrasting data suggest that efficiency gains from GenAI are not uniform, often redistributing development effort from content creation to verification and oversight (DeBellis et al., 2024; Harness, 2025; Stack Overflow, 2025c).
@@
Organizations, however, are adopting genAI at a rapid pace, and in the process escalating developer productivity expectations and job automations (HackerRank, 2025). Labor markets reflect these trends: the Stanford Digital Economy Lab reported entry-level cognitive jobs, disproportionately held by early-career developers, have declined by 13–20% (Brynjolfsson et al., 2025; Bernard, 2025). Other reports also show reduced headcounts because of AI automation (Cerullo, 2025; Leopold, 2025). Automation has another hidden cost on the workforce; it causes reduced situational awareness and decision-making skills, leading to actual (Kaber and Endsley, 1997) and perceived fear of deskilling (Crowston et al., 2025). To avoid such deskilling and to keep up with the rapid pace of AI development, developers have to continually adapt and reskill to stay relevant (Qiu et al., 2025; Necula, 2023).
@@
Burnout—characterized by a chronic state of exhaustion, emotional distance, cynicism, and diminished professional efficacy (Schaufeli and Bakker, 2004; Lazarus and Folkman, 1985; Anand et al., 2023)—is a foreseeable outcome in this situation.
@@
Burnout itself is not new to Software Engineering (SE); it has surfaced during prior disruptions, from the rise of agile (Tulili et al., 2023) to the COVID-19 pandemic (Cao et al., 2024). Past works, however, have not theorized the impact of genAI adoption on burnout. Instead, these studies have largely focused on developer productivity (e.g., speedups, throughput) (Houck et al., 2025; Becker et al., 2025), creativity (Jackson et al., 2024), or adoption intent (Russo, 2024; Johnson et al., 2023; Choudhuri et al., 2025), without researching its effects on developers’ well-being.
@@
We answer these questions with a concurrent embedded study (Greene, 2007; Morse, 2016) of a survey of professional developers ($N=442$), integrating quantitative modeling with qualitative evidence through the lens of the Job Demands–Resources (JD–R) model. The JD–R model is particularly well-suited for our context because it conceptualizes burnout as the outcome of an imbalance between job demands (e.g., organizational pressure, workload) and job resources (e.g., autonomy, learning resources), and has been applied in investigating burnout in other domains, such as the impact of COVID-19 (Cao et al., 2024).
@@
Our contributions are threefold:
  • We establish that GenAI adoption is linked to burnout through elevated organizational pressure and workload, while autonomy and learning resources, which are unevenly distributed, mitigate these effects.
  • We offer the first explanatory model of when GenAI acts as a demand amplifier versus a resource enabler in software development by extending the JD-R model to AI adoption.
  • We translate our findings into implications for workload design, workforce development, and team practices to help organizations enable equitable workload and resources for their employees.
@@ 2. Background
GenAI-powered tools are being rapidly integrated into software development workflows (Nguyen-Duc et al., 2025; Research, [n. d.]; GitHub, 2023), fueled by the promise of productivity gains. For example, developers have reported task speed-ups ranging from 21% to 55% when using GenAI tools (Peng et al., 2023; Paradis et al., 2025; Kalliamvakou, 2022). GitHub Copilot, for example, has been credited with cutting development lead times by more than half (Kalliamvakou, 2022). Houck et al. (2025) utilized the SPACE framework to demonstrate that AI significantly enhances productivity, particularly for routine tasks.
@@
These results provide a powerful narrative for organizations to actively encourage developers to integrate GenAI tools, so as to accelerate efficiency and remain competitive (Li et al., 2024). A growing body of research is investigating how GenAI can be effectively integrated into software development workflows. For example, Esposito et al. (2025) reviewed 46 studies from over 1,000 screened papers, focusing on how GenAI is applied in software architecture, including its potential to support design decisions. He et al. (2025) synthesized 71 studies on LLM-based multi-agent systems, outlining integration pathways across requirements engineering, code generation, quality assurance, and maintenance.
@@
Studies have also investigated the different human factors that influence GenAI adoption (Afroogh et al., 2024). For example, Huynh and Aichner (2025) emphasized that fairness, social presence, and emotions are determinants of trust in GenAI, which in turn shapes user attitudes, perceived performance, and adoption intentions. Choudhuri et al. (2025) demonstrated that developers’ trust and adoption intentions are shaped not only by system quality and functional value but also by cognitive diversity.
@@
But these positive narratives are not uniform, and there are perils to the GenAI hype (Kang et al., 2024; Stack Overflow, 2025a). A survey of 39,000 developers found only a 2.1% rise in overall productivity and a 3.4% improvement in code quality for every 25% increase in AI adoption, while software delivery performance actually declined by 7.2% (DeBellis et al., 2024). Becker et al. (2025) found that AI tools slowed developers down, increasing task completion time by 19%.
@@
Beyond efficiency, GenAI adoption introduces new job demands: on one hand, developers face organizational pressure to quickly adopt AI tools; on the other hand, once adopted, they are expected to work at a faster pace (Google / DORA / Google Cloud, 2025; Jaworski and Piotrkowski, 2023). At the same time, developers must invest substantial effort in validating, debugging, and securing AI-generated outputs, which often displaces the time saved on initial coding tasks (Harness, 2025; Stack Overflow, 2025c). However, amid this hype and intensified workload, developers report receiving insufficient training and organizational support (Mayer et al., 2025; Gao et al., 2025).
@@
Together, these emerging observations suggest that declining performance, rising organizational pressure, heavier workload, and insufficient resources are creating a growing strain for developers. Yet the well-being of developers impacted by these changes remains largely unexplored within the SE society.
@@
To address this gap, we turn to the construct of burnout, which captures how work-related stress translates into exhaustion, detachment, and diminished efficacy (Maslach et al., 2001). Burnout can manifest through multiple interrelated aspects of work life, such as feeling drained by demands, developing detached or negative attitudes toward work (Schaufeli and Bakker, 2004), and experiencing uncertainty about job security (Anand et al., 2023). Several theories could be applied to analyze burnout, including the Transactional Model of Stress (Lazarus and Folkman, 1985), the Conservation of Resources theory (Hobfoll, 1989), Social Cognitive Theory (Bandura, 1991), and Job Demands–Resources (JD–R) model (Demerouti et al., 2001; Bakker and Demerouti, 2017). We use the JD-R model as our analytical lens.
@@
The JD–R model (Demerouti et al., 2001; Bakker and Demerouti, 2017), developed initially as a general model of burnout (Demerouti et al., 2001), conceptualizes how job demands (e.g., workload, pressure, uncertainty) and job resources (e.g., autonomy, learning opportunities) jointly shape well-being. Its flexibility has made it a widely adopted framework for investigating well-being across diverse work contexts, including software development (Bakker and Demerouti, 2017; Trinkenreich et al., 2024b; Kim et al., 2024).
@@ 3. Theory Development
Job Demands refer to the physical, social, or organizational aspects of work that require sustained cognitive effort and are associated with physiological and psychological costs, such as exhaustion (Demerouti et al., 2001; Bakker and Demerouti, 2017). Foundational JD–R studies (Balducci et al., 2011; Bakker and Demerouti, 2007) conceptualize Job Demands as including key factors like (1) Organizational Pressure and (2) Workload.
@@
Organizational pressure. In our context, leadership-driven AI adoption mandates can lead to ineffective AI integration when organizational readiness does not align with leadership expectations (Mayer et al., 2025), which can cause structural strain on teams who are pushed to adopt AI with ineffective results. Even when there is organizational readiness, developers’ workload can increase if they are assigned more tasks or expected to complete tasks at a faster pace by management because of AI adoption. Workload can also increase if developers have to spend additional time debugging or fixing AI errors (Harness, 2025). Past work has shown that these complementary factors, organizational pressure and workload, are significantly associated with elevated work stress (Roșca et al., 2021; Scholze and Hecker, 2024). We therefore hypothesize: H1: Job demands related to GenAI adoption are associated with developer burnout.
@@
Job Resources refer to the physical, social, or organizational aspects of work that help individuals achieve work goals, reduce job demands, or stimulate personal growth and development (Demerouti et al., 2001; Bakker and Demerouti, 2017). It includes the two formative components: (1) Autonomy and (2) Organizational support as its constituents, both of which have been shown to buffer the adverse effects of job demands on burnout (Bakker and Demerouti, 2007).
@@
Autonomy in our context reflects the degree of control developers retain over how they integrate AI tools into their work, and is a well-known factor in improving employee well-being, motivation, and performance when navigating changing work environments (Ravn et al., 2022). Organizational support, such as providing learning resources, can enable developers to maximize the value of AI (Houck et al., 2025). Other, non-AI-related research has consistently shown that autonomy and organizational support are important aspects of job resources to reduce burnout and foster engagement in technology-driven work environments (Scholze and Hecker, 2024; Roșca et al., 2021). Similarly, reviews of JD–R applications in leadership contexts emphasize autonomy as a central resource for mitigating strain and enhancing performance (Tummers and Bakker, 2021; Bakker et al., 2004). Building on these insights, we hypothesize that: H2: Job resources related to GenAI adoption are associated with developer burnout.
@@
Burnout is a state of exhaustion, cynicism, and reduced efficiency caused by prolonged stressors, including work pressures and intrinsic factors such as mental stress or job insecurity (Maslach et al., 2001). Organizational mandates for GenAI adoption can amplify job insecurity when routine development tasks are advertised as being automated or when management signals potential workforce reductions (Molino et al., 2020). Additionally, constant news of AI adoption as well as the breakneck speed of AI advancement can induce fear of deskilling (Crowston et al., 2025). However, burnout, a complex social construct, can be mitigated when an individual is driven and excited by new advancements (in GenAI). Research shows that favorable technology beliefs (e.g., usefulness and trust) can reduce strain and mitigate negative outcomes (Davis, 1989; Venkatesh et al., 2012). Therefore, we hypothesize that H3: Favorable perceptions of AI are associated with developer burnout.
@@
Developer Characteristics. The effects of job demands and access to resources are not uniform across an organization and are dependent on factors such as seniority, role, or organizational context (Demerouti et al., 2001). To examine whether these characteristics influence different constructs of the JD–R model, we hypothesize: H4: Developer characteristics are associated with variation in: Burnout (H4a); Organizational Pressure (H4b); Workload (H4c); Autonomy (H4d); and Learning Resources (H4e).
@@
In the upcoming sections, we first describe the survey design and data collection process in Section 4. In Section 5 we evaluate H1–H3 using Partial Least Squares–Structural Modeling (PLS-SEM) to understand the relationships among job demands, job resources, AI perceptions, and burnout (RQ1). Then, in Section 6, we assess H4a–H4e through regression analyses of developers’ characteristics (role, industry seniority, organization size) to assess variation in demands, resources, and burnout (RQ2). We then complement these quantitative analyses with qualitative reporting of open-ended responses to contextualize developers’ experiences with GenAI adoption.
@@ 4. Data Collection: Survey
We collected data to test the above hypotheses through a survey of software professionals. Three researchers with experience in survey methodology and software engineering research collaboratively developed the survey instrument questions over the course of one month using the JD–R model as a guide (Demerouti et al., 2001). The constructs in the JD–R model are defined as latent constructs: constructs that cannot be directly measured or observed, and instead are measured through a set of indicators (survey items). We adapted existing survey questions to measure the different constructs.
@@
Survey design. The survey began with the university’s Institutional Review Board (IRB) approved consent form, followed by five sections comprising a total of 15 questions. See the complete questionnaire in the supplementary document Section 1 (Anonym, 2025).
@@
The first survey section collected participant characteristics, including gender, professional background (primary role in software development, years of industry experience, and seniority level); organizational context (organization size), AI-usage experience (frequency of AI use in software development work, adapted from (Research, [n. d.])), and attitude towards AI adoption measured by agreement with the statement: “I feel positive about AI usage”.
@@
The following section, related to Job Demands, asked questions about organizational pressure and workload resulting from AI adoption. We measured organizational pressure by asking participants to rate 4 statements on a 5-point Likert scale (1 = strongly disagree, 5 = strongly agree; N/A if not applicable). These questions were adapted from work on technology adoption and job design (Brown et al., 2002; Morgeson and Humphrey, 2006), which showed that organizational mandates, collaborative expectations, and heightened responsibility can generate pressure. Some sample questions are: “I must also use AI to get my work done as my co-workers rely on AI outputs”; “My performance evaluations depend on how effectively I use AI tools”. Workload questions also used the same 5-point Likert scale and asked participants to rate 4 statements. These statements were adapted from Cao et al. (2024) work that used JD–R to examine the impact of COVID-19 on workload. Two example questions are: “I often have to do more work than I can do well”; “My job requires me to work harder”.
@@
The third survey section was about Job Resources and included questions related to autonomy and learning resources. We measured autonomy through questions, adapted from Clausen et al. (2019), about participants’ perceived discretion and decision-making power in their development work, using a 5-point Likert scale (from “To a very small extent” to “To a very large extent”). Some example questions are: “Do you have any influence on how you carry out your work tasks?”; “Do you have sufficient authority to deal with the responsibilities you have in your work?”. For measuring availability of resources for learning AI, we asked participants to rate the statement: “My organization gives me the time, training, and resources I need to master AI” on a 5-point Likert scale (1 = strongly disagree, 5 = strongly agree; N/A if not applicable). The survey then followed with an open-ended question about the AI training provided by participants’ employers, adapted from (Oliver Wyman Forum, 2024).
@@
The last survey section measured burnout through participants’ perceptions through a set of 4 statements which participants rated on a 5-point Likert scale (1 = strongly disagree, 5 = strongly agree; N/A if not applicable). These questions were adapted from (Trinkenreich et al., 2024b) to capture established dimensions of burnout: workload manageability and exhaustion, loss of interest as an indicator of cynicism, and insecurity about one’s future as a reflection of reduced efficacy. Sample questions include: “I feel mentally and physically exhausted from work”; “I feel secure in my job and confident about my future in this company”.
@@
Sandbox and Pilot Survey. Before distributing the survey, we sandboxed the survey with five participants who had research experience in software engineering. The survey was iteratively refined through the sandbox sessions until no further concerns remained. We piloted the survey with two collaborators: one from a multinational technology corporation and the other from a U.S. national laboratory. Both found the survey straightforward, easy to understand, and relevant.
Participant Recruitment. To ensure broad coverage across technical backgrounds, seniority levels, organization sizes, and project domains, we recruited software developers from 56 Open Source Software (OSS) communities spanning diverse domains, including organizational repositories (e.g., Microsoft, Google, Netflix, Red Hat, IBM) as well as widely used open-source infrastructure, cloud, and AI projects (e.g., Kubernetes, Hugging Face). We also reached out to data science–oriented communities (e.g., Python, TensorFlow) and professional groups. We intentionally used a broad recruitment strategy to reflect the evolving nature of software development, which increasingly involves not only traditional software engineers but also data scientists, machine learning practitioners, and AI developers working together in collaborative software development environments (Feng et al., 2025b).
In accordance with ethical data collection guidelines, we sent email invitations detailing the survey’s purpose, along with a consent form that contained information on the survey’s voluntary nature, potential risks, and data protection measures. All participants provided informed consent, and responses were anonymized in compliance with the General Data Protection Regulation (GDPR) and our IRB approval. The survey was available for two weeks. This recruitment approach aligns with prior SE research (Feng et al., 2022, 2025a, 2025b).

### v2

# From Gains to Strains: Modeling Developer Burnout with GenAI Adoption

@@ Abstract.
journalyear: 2026copyright: ccconference: 2026 IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 2026 IEEE/ACM 48th International Conference on Software Engineering (ICSE-SEIS ’26), April 12–18, 2026, Rio de Janeiro, Brazildoi: 10.1145/3786581.3786934isbn: 979-8-4007-2424-4/2026/04ccs: Human-centered computing Empirical studies in collaborative and social computingccs: Human-centered computing
@@ 1. Introduction
We are now at a pivotal shift in how Generative AI (GenAI) is transforming how we work and live. Since late 2022, GenAI has progressed from experimental pilots to mainstream adoption, reshaping how code is written, reviewed, and maintained (nguyen2025generative; oliverwyman2024ai). Individual adoption is rapid—84% of developers in the 2025 Stack Overflow survey reported using or planning to use GenAI tools (StackOverflow2025_AItools). Organizations are not far behind. Motivated by promises of higher productivity, they are heavily investing in GenAI (GitHub2025_CopilotImpact).
@@
Some studies support these investments. For example, GitHub reported that developers completed coding tasks 55.8% faster (GitHub2025_CopilotImpact). Yet, others report increased costs. For instance, the 2025 State of Software Delivery Report by Harness (harness2025) reported 67% of developers spent more time debugging AI-generated code and 68% spent more time fixing AI-created security issues. Similarly, becker2025measuring reported an overall 19% productivity loss in their observation study. These contrasting data suggest that efficiency gains from GenAI are not uniform, often redistributing development effort from content creation to verification and oversight (dora2024; harness2025; StackOverflow2025_AItools).
@@
Organizations, however, are adopting genAI at a rapid pace, and in the process escalating developer productivity expectations and job automations (HackerRank2025_PressureToDeliver). Labor markets reflect these trends: the Stanford Digital Economy Lab reported entry-level cognitive jobs, disproportionately held by early-career developers, have declined by 13–20% (brynjolfsson2025canaries; bernard2025tech). Other reports also show reduced headcounts because of AI automation (Cerullo2025_CBSjobsLayoffs; WEF2025_AIJobsLabourDay). Automation has another hidden cost on the workforce; it causes reduced situational awareness and decision-making skills, leading to actual (kaber1997out) and perceived fear of deskilling (crowston2025deskilling). To avoid such deskilling and to keep up with the rapid pace of AI development, developers have to continually adapt and reskill to stay relevant (qiu2025today; necula2023artificial).
@@
Burnout—characterized by a chronic state of exhaustion, emotional distance, cynicism, and diminished professional efficacy (schaufeli2004job; lazarus1985stress; anand2023effect)—is a foreseeable outcome in this situation.
@@
Burnout itself is not new to Software Engineering (SE); it has surfaced during prior disruptions, from the rise of agile (tulili2023burnout) to the COVID-19 pandemic (cao2024jd). Past works, however, have not theorized the impact of genAI adoption on burnout. Instead, these studies have largely focused on developer productivity (e.g., speedups, throughput) (houck2025space; becker2025measuring), creativity (jackson2024creativity), or adoption intent (russo2024navigating; johnson2023make; choudhuri2025needs), without researching its effects on developers’ well-being.
@@
We answer these questions with a concurrent embedded study (greene2007mixed; morse2016mixed) of a survey of professional developers ($N=442$), integrating quantitative modeling with qualitative evidence through the lens of the Job Demands–Resources (JD–R) model. The JD–R model is particularly well-suited for our context because it conceptualizes burnout as the outcome of an imbalance between job demands (e.g., organizational pressure, workload) and job resources (e.g., autonomy, learning resources), and has been applied in investigating burnout in other domains, such as the impact of COVID-19 (cao2024jd).
@@
Our contributions are threefold: (1) We establish that GenAI adoption is linked to burnout through elevated organizational pressure and workload, while autonomy and learning resources, which are unevenly distributed, mitigate these effects; (2) We offer the first explanatory model of when GenAI acts as a demand amplifier versus a resource enabler in software development by extending the JD–R model to AI adoption; (3) We translate our findings into implications for workload design, workforce development, and team practices to help organizations enable equitable workload and resources for their employees.
@@ 2. Background
GenAI-powered tools are being rapidly integrated into software development workflows (nguyen2025generative; usage2025; GitHub2025_CopilotImpact), fueled by the promise of productivity gains (qiao2025comprehension; shihab2025effects). For example, developers have reported task speed-ups ranging from 21% to 55% when using GenAI tools (peng2023impact; paradis2025much; kalliamvakou2022copilot). GitHub Copilot, for example, has been credited with cutting development lead times by more than half (kalliamvakou2022copilot). houck2025space utilized the SPACE framework to demonstrate that AI significantly enhances productivity, particularly for routine tasks.
@@
These results provide a powerful narrative for organizations to actively encourage developers to integrate GenAI tools, so as to accelerate efficiency and remain competitive (li2024ai). A growing body of research is investigating how GenAI can be effectively integrated into software development workflows. For example, esposito2025generative reviewed 46 studies from over 1,000 screened papers, focusing on how GenAI is applied in software architecture, including its potential to support design decisions. he2025llm synthesized 71 studies on LLM-based multi-agent systems, outlining integration pathways across requirements engineering, code generation, quality assurance, and maintenance.
@@
Studies have also investigated the different human factors that influence GenAI adoption (afroogh2024trust). For example, huynh2025generative emphasized that fairness, social presence, and emotions are determinants of trust in GenAI, which in turn shapes user attitudes, perceived performance, and adoption intentions. choudhuri2025needs demonstrated that developers’ trust and adoption intentions are shaped not only by system quality and functional value but also by cognitive diversity.
@@
But these positive narratives are not uniform, and there are perils to the GenAI hype (kang2024quantitative; StackOverflow2025_AItools). A survey of 39,000 developers found only a 2.1% rise in overall productivity and a 3.4% improvement in code quality for every 25% increase in AI adoption, while software delivery performance actually declined by 7.2% (dora2024). becker2025measuring found that AI tools slowed developers down, increasing task completion time by 19%.
@@
Beyond efficiency, GenAI adoption introduces new job demands: on one hand, developers face organizational pressure to quickly adopt AI tools; on the other hand, once adopted, they are expected to work at a faster pace (google2025state; jaworski2023study). At the same time, developers must invest substantial effort in validating, debugging, and securing AI-generated outputs, which often displaces the time saved on initial coding tasks (harness2025; StackOverflow2025_AItools). However, amid this hype and intensified workload, developers report receiving insufficient training and organizational support (mayer2025superagency; gao2025ai).
@@
Together, these emerging observations suggest that declining performance, rising organizational pressure, heavier workload, and insufficient resources are creating a growing strain for developers. Yet, the well-being of developers impacted by these changes remains largely unexplored within the SE society.
@@
To address this gap, we turn to the construct of burnout, which captures how work-related stress translates into exhaustion, detachment, and diminished efficacy (maslach2001job). Burnout can manifest through multiple interrelated aspects of work life, such as feeling drained by demands, developing detached or negative attitudes toward work (schaufeli2004job), and experiencing uncertainty about job security (anand2023effect). Several theories could be applied to analyze burnout, including the Transactional Model of Stress (lazarus1985stress), the Conservation of Resources theory (hobfoll1989conservation), Social Cognitive Theory (bandura1991social), and Job Demands–Resources (JD–R) model (demerouti2001job; bakker2017job). We use the JD–R model as our analytical lens.
@@
The JD–R model (demerouti2001job; bakker2017job), developed initially as a general model of burnout (demerouti2001job), conceptualizes how job demands (e.g., workload, pressure, uncertainty) and job resources (e.g., autonomy, learning opportunities) jointly shape well-being. Its flexibility has made it a widely adopted framework for investigating well-being across diverse work contexts, including software development (bakker2017job; trinkenreich2024predicting).
@@ 3. Theory Development
Job Demands refer to the physical, social, or organizational aspects of work that require sustained cognitive effort and are associated with physiological and psychological costs, such as exhaustion (demerouti2001job; bakker2017job). Foundational JD–R studies (balducci2011job; bakker2007job) conceptualize Job Demands as including key factors like (1) Organizational Pressure and (2) Workload.
@@
Organizational pressure. In our context, leadership-driven AI adoption mandates can lead to ineffective AI integration when organizational readiness does not align with leadership expectations (mayer2025superagency), which can cause structural strain on teams who are pushed to adopt AI with ineffective results. Even when there is organizational readiness, developers’ workload can increase if they are assigned more tasks or expected to complete tasks at a faster pace by management because of AI adoption. Workload can also increase if developers have to spend additional time debugging or fixing AI errors (harness2025). Past work has shown that these complementary factors, organizational pressure and workload, are significantly associated with elevated work stress (rosca2021job; scholze2024job). We therefore hypothesize: H1: Job demands related to GenAI adoption are associated with developer burnout.
@@
Job Resources refer to the physical, social, or organizational aspects of work that help individuals achieve work goals, reduce job demands, or stimulate personal growth and development (demerouti2001job; bakker2017job). It includes the two formative components: (1) Autonomy and (2) Organizational support as its constituents, both of which have been shown to buffer the adverse effects of job demands on burnout (bakker2007job).
@@
Autonomy in our context reflects the degree of control developers retain over how they integrate AI tools into their work, and is a well-known factor in improving employee well-being, motivation, and performance when navigating changing work environments (ravn2022team). Organizational support, such as providing learning resources, can enable developers to maximize the value of AI (houck2025space). Other, non-AI-related research has consistently shown that autonomy and organizational support are important aspects of job resources to reduce burnout and foster engagement in technology-driven work environments (scholze2024job; rosca2021job). Similarly, reviews of JD–R applications in leadership contexts emphasize autonomy as a central resource for mitigating strain and enhancing performance (tummers2021leadership; bakker2004using). Building on these insights, we hypothesize that: H2: Job resources related to GenAI adoption are associated with developer burnout.
@@
Burnout is a state of exhaustion, cynicism, and reduced efficiency caused by prolonged stressors, including work pressures and intrinsic factors such as mental stress or job insecurity (maslach2001job). Organizational mandates for GenAI adoption can amplify job insecurity when routine development tasks are advertised as being automated or when management signals potential workforce reductions (molino2020promotion). Additionally, constant news of AI adoption as well as the breakneck speed of AI advancement can induce fear of deskilling (crowston2025deskilling). However, burnout, a complex social construct, can be mitigated when an individual is driven and excited by new advancements (in GenAI). Research shows that favorable technology beliefs (e.g., usefulness and trust) can reduce strain and mitigate negative outcomes (davis1989perceived; venkatesh2012consumer). Therefore, we hypothesize that H3: Favorable perceptions of AI are associated with developer burnout.
@@
Developer Characteristics. The effects of job demands and access to resources are not uniform across an organization and are dependent on factors such as seniority, role, or organizational context (demerouti2001job). To examine whether these characteristics influence different constructs of the JD–R model, we hypothesize: H4: Developer characteristics are associated with variation in: Burnout (H4a); Organizational Pressure (H4b); Workload (H4c); Autonomy (H4d); and Learning Resources (H4e).
@@
In the upcoming sections, we first describe the survey design and data collection process in Section 4. In Section 5, we evaluate H1–H3 using Partial Least Squares–Structural Modeling (PLS-SEM) to understand the relationships among job demands, job resources, AI perceptions, and burnout (RQ1). Then, in Section 6, we assess H4a–H4e through regression analyses of developers’ characteristics (role, industry seniority, organization size) to assess variation in demands, resources, and burnout (RQ2). We then complement these quantitative analyses with qualitative reporting of open-ended responses to contextualize developers’ experiences with GenAI adoption.
@@ 4. Data Collection: Survey
We collected data to test the above hypotheses through a survey of software professionals. Three researchers with experience in survey methodology and software engineering research collaboratively developed the survey instrument questions over the course of one month using the JD–R model as a guide (demerouti2001job). The constructs in the JD–R model are defined as latent constructs: constructs that cannot be directly measured or observed, and instead are measured through a set of indicators (survey items). We adapted existing survey questions to measure the different constructs.
@@
Survey design. The survey began with the university’s Institutional Review Board (IRB) approved consent form, followed by five sections comprising a total of 15 questions. See the complete questionnaire in the supplementary document Section 1 (supply).
@@
The first survey section collected participant characteristics, including gender, professional background (primary role in software development, years of industry experience, and seniority level); organizational context (organization size), AI-usage experience (frequency of AI use in software development work, adapted from (usage2025)), and attitude towards AI adoption measured by agreement with the statement: “I feel positive about AI usage”.
@@
The following section, related to Job Demands, asked questions about organizational pressure and workload resulting from AI adoption. We measured organizational pressure by asking participants to rate 4 statements on a 5-point Likert scale (1 = strongly disagree, 5 = strongly agree; N/A if not applicable). These questions were adapted from work on technology adoption and job design (brown2002really; morgeson2006work), which showed that organizational mandates, collaborative expectations, and heightened responsibility can generate pressure. Some sample questions are: “I must also use AI to get my work done as my co-workers rely on AI outputs”; “My performance evaluations depend on how effectively I use AI tools”. Workload questions also used the same 5-point Likert scale and asked participants to rate 4 statements. These statements were adapted from cao2024jd work that used JD–R to examine the impact of COVID-19 on workload. Two example questions are: “I often have to do more work than I can do well”; “My job requires me to work harder”.
@@
The third survey section was about Job Resources and included questions related to autonomy and learning resources. We measured autonomy through questions, adapted from clausen2019danish, about participants’ perceived discretion and decision-making power in their development work, using a 5-point Likert scale (from “To a very small extent” to “To a very large extent”; N/A if not applicable). Some example questions are: “Do you have any influence on how you carry out your work tasks?”; “Do you have sufficient authority to deal with the responsibilities you have in your work?”. For measuring availability of resources for learning AI, we asked participants to rate the statement: “My organization gives me the time, training, and resources I need to master AI” on a 5-point Likert scale (1 = strongly disagree, 5 = strongly agree; N/A if not applicable). The survey then followed with an open-ended question about the AI training provided by participants’ employers, adapted from (oliverwyman2024ai).
@@
The last survey section measured burnout through participants’ perceptions through a set of 4 statements which participants rated on a 5-point Likert scale (1 = strongly disagree, 5 = strongly agree; N/A if not applicable). These questions were adapted from trinkenreich2024predicting to capture established dimensions of burnout: workload manageability and exhaustion, loss of interest as an indicator of cynicism, and insecurity about one’s future as a reflection of reduced efficacy. Sample questions include: “I feel mentally and physically exhausted from work”; “I feel secure in my job and confident about my future in this company”.
@@
Sandbox and Pilot Survey. Before distributing the survey, we sandboxed the survey with five participants who had research experience in SE. The survey was iteratively refined through the sandbox sessions until no further concerns remained. We piloted the survey with two collaborators: one from a multinational technology corporation and the other from a U.S. national laboratory. Both found the survey straightforward, easy to understand, and relevant.
Participant Recruitment. To ensure broad coverage across technical backgrounds, seniority levels, organization sizes, and project domains, we recruited software developers from 56 Open Source Software (OSS) communities spanning diverse domains, including organizational repositories (e.g., Microsoft, Google, Netflix, Red Hat, IBM) as well as widely used open-source infrastructure, cloud, and AI projects (e.g., Kubernetes, Hugging Face). We also reached out to data science-oriented communities (e.g., Python, TensorFlow) and professional groups. We intentionally used a broad recruitment strategy to reflect the evolving nature of software development, which increasingly involves not only traditional software engineers but also data scientists, machine learning practitioners, and AI developers working together in collaborative software development environments (feng2025domains).
In accordance with ethical data collection guidelines, we sent email invitations detailing the survey’s purpose, along with a consent form that contained information on the survey’s voluntary nature, potential risks, and data protection measures. All participants provided informed consent, and responses were anonymized in compliance with the General Data Protection Regulation (GDPR) and our IRB approval. The survey was available for two weeks. This recruitment approach aligns with prior SE research (feng2022case; feng2025multifaceted; feng2025domains).